i need help with solving rational numbers with fractions.
(2x-3) = x-4 - x+1
4 2 4 Help!!!

Answer by checkley77(12844)

(2x-3) = x-4 - x+1
4 2 4
TO GET RID OF ALL THE DENOMINATORS BY MULTIPLYING ALL TERMS BY 4.
2X-3=2(X-4)-X-1
2X-3=2X-8-X-1
2X-2X+X=-8-1+3
X=-6 ANSWER.
PROOF:
(2*-6-3)/4=(-6-4)/2-(-6+1)/4
(-12-3)/4=-10/2-(-5/4)
-15/4=-5+5/4
-15/4=-20/4+5/4
-15/4=-15/4